OpenStudy (ivitt):

50n(678*n)+78*999-900=9000 Halp Please

OpenStudy (quickstudent):

Are you trying to solve for n? It would come out as an irrational number.

OpenStudy (sshayer):

\[50 n(678n)+78(1000-1)-900=9000\] \[50n(678n)+78,000-78-900=9000 \] 50n(678n)=9000+900+78-78,000 50 n (678 n)=9,978-78,000=-68022 50*678 n^2=-68022 n is an imaginary number.